# Time-asymptotic self-similarity of the damped compressible Euler equations in parabolic scaling variables

### v1

@@ 5.2 Reformulation in parabolic scaling variables
As in the one-dimensional case, to study the long-time behavior of solutions to (5.1), we consider the problem in parabolic scaling variables. We define $\tau:=\log(1{+}t)$ and $y:=\frac{x}{\sqrt{1{+}t}}$ and set $\rho(\tau,y):=\widetilde{\rho}(t,x)$ and $n(\tau,y):=\mathrm{e}^{d\tau/2}m(\tau,y)=(1{+}t)^{1/2}\,\widetilde{m}(t,x)$. Observe that in the multi-dimensional framework, we scale the magnitude of $\widetilde{m}$ by $\mathrm{e}^{d\tau/2}$ instead of $\mathrm{e}^{\tau/2}$. Then (5.1) is equivalent to
@@
$n_{\tau}$ $=\frac{y}{2}\cdot\nabla n+\frac{1}{2}n-\operatorname{div}\Big{(}\frac{n\otimes n}{\rho}\Big{)}-\mathrm{e}^{d\tau}\big{(}\nabla\mathfrak{p}(\rho)_{y}+\alpha n\big{)}.$
@@
Let $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d};{\mathbb{R}}^{d})$ such that ${\mathcal{E}}(\rho_{0},n_{0}\,|\,\hat{\rho},0)<\infty$. We call a pair $(\rho,n,E)\in L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d};{\mathbb{R}}^{d})\times\mathrm{BV}([0,\infty))$ an energy-variational solution to (5.7) if there is $\rho_{\infty}\in C^{1}([0,\infty)\times{\mathbb{R}}^{d})$ with (5.8) such that ${\mathcal{E}}(\tau;\rho,n\,|\,\rho_{\infty},0)\leq E$ a.e. in $(0,\infty)$, and if there exists a $1$-homogeneous regularity weight $\mathcal{K}\colon C_{c}^{1}({\mathbb{R}}^{d};{\mathbb{R}}^{d})\to[0,\infty)$ such that
@@ 5.3 Relative entropy inequality and long-time behavior
where $\overline{R}:=\frac{\overline{n}}{\overline{\rho}}\cdot\overline{R}_{1}-\overline{R}_{2}$ and
@@
$\overline{R}_{2}$ $:=\overline{n}_{\tau}-\frac{y}{2}\cdot\nabla\overline{n}-\frac{1}{2}\overline{n}+\operatorname{div}\Big{(}\frac{\overline{n}\otimes\overline{n}}{\overline{\rho}}\Big{)}+\mathrm{e}^{d\tau}\big{(}\nabla\mathfrak{p}(\overline{\rho})+\alpha\overline{n}\big{)},$
@@
    $\overline{\rho}-\hat{\rho}$ and $\overline{n}$ have compact support in space, or if $\hat{\rho}\equiv 0$, then it holds $(\overline{\rho},\overline{n})=(0,0)$ and $\mathfrak{p}$ satisfy (2.14), or
@@
$\qquad\qquad-\int_{0}^{\infty}\!\!\varphi\int_{{\mathbb{R}}^{d}}\psi_{\tau}\,n+\nabla\psi:\Big{(}-n\otimes\frac{y}{2}+\frac{n\otimes n}{\rho}+\mathrm{e}^{d\tau}\mathfrak{p}(\rho){\mathbb{I}}\Big{)}-\psi\cdot\mathrm{e}^{d\tau}\alpha n\,\mathrm{d}y\,\mathrm{d}\tau$
$\qquad\qquad\qquad\leq\varphi(0)\Big{(}E(0)+\int_{{\mathbb{R}}^{d}}\phi(0,y)\rho(0,y)+\psi(0,y)\cdot n(0,y)\,\mathrm{d}y\Big{)}$
@@
for all $\phi,\psi\in C_{c}^{1}([0,\infty)\times{\mathbb{R}})$ and all $\varphi\in C_{c}^{1}([0,\infty))$ with $\varphi\geq 0$. Observe that we can also consider test functions that are merely Lipschitz continuous, and that $\phi$ and $\psi$ need not have compact support in time due to the presence of $\varphi$. We consider $\phi(\tau,y)=\frac{1}{2}\mathrm{e}^{-\tau}\big{(}\frac{\overline{n}(\tau,y)}{\overline{\rho}(\tau,y)}\big{)}^{2}+h^{\prime}(\rho_{\infty}(\tau,y))-h^{\prime}(\overline{\rho}(\tau,y))$ and $\psi(\tau,y)=-\mathrm{e}^{-\tau}\frac{\overline{n}(\tau,y)}{\overline{\rho}(\tau,y)}$. Then a direct calculation that uses the same identities as in the proof of Lemma 3.12 leads to to (5.16) in case (a). In case (b), we can subsequently use the approximation argument from the proof of Theorem 3.14, using the coercivity estimate from Lemma 3.4, to extend the class of admissible functions $(\overline{\rho},\overline{n})$ as asserted. ∎
@@
Assumption (A) is clearly satisfied if $\hat{\rho}$ is constant, which yields $\rho^{*}\equiv\hat{\rho}$. Moreover, if $\mathfrak{p}$ is of the form (1.2), then one can construct a smooth self-similar solution to the porous medium equation $\rho_{t}=\frac{1}{\alpha}\Delta\mathfrak{p}(\rho)$, which leads to a similarity profile $\rho^{*}$ satisfying (5.17), for instance, by following the ideas of [Váz07, Thm. 16.2]. However, the decay properties of $\rho^{*}$ are less clear, and in multiple dimensions, they may depend on the angular profile $\hat{\rho}$ and assumption (A) may be a proper restriction.
@@
    Let $\alpha\geq 0$, let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$ in $(0,\infty)$, let $\rho^{*}=\hat{\rho}\geq 0$ be constant and $n^{*}=0$, and if $\rho^{*}=0$, let $\mathfrak{p}$ satisfy (2.14), or
@@
$\mu$ $:=\Big{\|}\frac{|R^{*}|}{2k(\rho^{*})^{\gamma}}+\frac{3|R^{*}|}{2\rho^{*}}\Big{\|}_{L^{\infty}({\mathbb{R}})},\qquad\quad K:=\int_{\mathbb{R}}|R^{*}|\,\mathrm{d}y,$
@@
where we used ${\mathcal{E}}(\tau;\rho,n\,|\,\rho^{*},n^{*})\leq\mathcal{R}$ for the first term in the integral on the right-hand side. Invoking now Gronwall’s inequality from Lemma 3.15 as in the proof of Theorem 2.3, we conclude (5.18) in this case.
@@
Theorem 5.10 yields convergence of energy-variational solutions towards the similarity profile under the same flatness condition as for weak solutions, see Theorem 2.3. In a similar way as in Section 4, we could also deduce convergence of the time-avereaged kinetic energy. In Theorem 5.10, we further deduce the convergence of the energy defect to zero, which means that energy-variational solutions converge to weak solutions in a certain sense. Observe that in the case $\rho_{+}=\rho_{-}$ the result holds for any regularity weight $\mathcal{K}$, while in the case $\rho_{+}\neq\rho_{-}$ we require the specific choice (5.6). Clearly, one could also take a regularity weight that is larger than (5.6), but this would lead to a more restrictive flatness condition.

### v2

@@ 5.2 Reformulation in parabolic scaling variables
As in the one-dimensional case, to study the long-time behavior of solutions to (5.1), we consider the problem in parabolic scaling variables. We define $\tau:=\log(1{+}t)$ and $y:=\frac{x}{\sqrt{1{+}t}}$ and set $\rho(\tau,y):=\widetilde{\rho}(t,x)$ and $n(\tau,y):=\mathrm{e}^{d\tau/2}m(\tau,y)=(1{+}t)^{d/2}\,\widetilde{m}(t,x)$. Observe that in the multi-dimensional framework, we scale the magnitude of $\widetilde{m}$ by $\mathrm{e}^{d\tau/2}$ instead of $\mathrm{e}^{\tau/2}$. Then (5.1) is equivalent to
@@
$n_{\tau}$ $=\frac{y}{2}\cdot\nabla n+\frac{d}{2}n-\operatorname{div}\Big{(}\frac{n\otimes n}{\rho}\Big{)}-\mathrm{e}^{d\tau}\big{(}\nabla\mathfrak{p}(\rho)_{y}+\alpha n\big{)}.$
@@
Let $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d};{\mathbb{R}}^{d})$ such that ${\mathcal{E}}(\rho_{0},n_{0}\,|\,\hat{\rho},0)<\infty$. We call a pair $(\rho,n,E)\in L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d};{\mathbb{R}}^{d})\times\mathrm{BV}([0,\infty))$ an energy-variational solution to (5.7) if there is $\rho_{\infty}\in C^{1}([0,\infty)\times{\mathbb{R}}^{d})$ with (5.8) such that ${\mathcal{E}}(\tau;\rho,n\,|\,\rho_{\infty},0)\leq E$ a.e. in (0,)(0$\infty$)( 0 , ∞ ), and if there exists a 111-homogeneous regularity weight $\mathcal{K}\colon C_{c}^{1}({\mathbb{R}}^{d};{\mathbb{R}}^{d})\to[0,\infty)$ such that
@@ 5.3 Relative entropy inequality and long-time behavior
where $\overline{R}:=\frac{\overline{n}}{\overline{\rho}}\cdot\overline{R}_{1}-\overline{R}_{2}$ with
@@
$\overline{R}_{2}$ :$=\overline{n}_{\tau}-\frac{y}{2}\cdot\nabla\overline{n}-\frac{d}{2}\overline{n}+\operatorname{div}\Big{(}\frac{\overline{n}\otimes\overline{n}}{\overline{\rho}}\Big{)}+\mathrm{e}^{d\tau}\big{(}\nabla\mathfrak{p}(\overline{\rho})+\alpha\overline{n}\big{)},$
@@
    $\overline{\rho}-\hat{\rho}$ and $\overline{n}$ have compact support in space, or if $\hat{\rho}\equiv 0$, then it holds $(\overline{\rho},\overline{n})=(0,0)$ and $\mathfrak{p}$ satisfies (2.14), or
@@
$\qquad\qquad-\int_{0}^{\infty}\!\!\varphi\int_{{\mathbb{R}}^{d}}\psi_{\tau}\cdot n+\nabla\psi:\Big{(}-n\otimes\frac{y}{2}+\frac{n\otimes n}{\rho}+\mathrm{e}^{d\tau}\mathfrak{p}(\rho){\mathbb{I}}\Big{)}-\psi\cdot\mathrm{e}^{d\tau}\alpha n\,\mathrm{d}y\,\mathrm{d}\tau$
$\qquad\qquad\qquad\leq\varphi(0)\Big{(}E(0)+\int_{{\mathbb{R}}^{d}}\phi(0,y)\rho_{0}(y)+\psi(0,y)\cdot n_{0}(y)\,\mathrm{d}y\Big{)}$
@@
for all $\phi,\psi\in C^{1}([0,\infty)\times{\mathbb{R}})$ and all $\varphi\in C_{c}^{1}([0,\infty))$ with $\varphi\geq 0$. Observe that $\phi$ and $\psi$ need not have compact support in time due to the presence of $\varphi$. We consider $\phi(\tau,y)=\frac{1}{2}\mathrm{e}^{-\tau}\big{(}\frac{\overline{n}(\tau,y)}{\overline{\rho}(\tau,y)}\big{)}^{2}+h^{\prime}(\rho_{\infty}(\tau,y))-h^{\prime}(\overline{\rho}(\tau,y))$ and $\psi(\tau,y)=-\mathrm{e}^{-\tau}\frac{\overline{n}(\tau,y)}{\overline{\rho}(\tau,y)}$. Then a direct calculation that uses the same identities as in the proof of Lemma 3.12 leads to (5.16) in case (a). In case (b), we can subsequently use the approximation argument from the proof of Theorem 3.14, using the coercivity estimate from Lemma 3.4, to extend the class of admissible functions $(\overline{\rho},\overline{n})$ as asserted. ∎
@@
Assumption (A) is clearly satisfied if $\hat{\rho}$ is constant, which yields $\rho^{*}\equiv\hat{\rho}$. Moreover, if $\mathfrak{p}$ is of the form (1.2), then one can construct a smooth self-similar solution to the porous medium equation $\widetilde{\rho}_{t}=\frac{1}{\alpha}\Delta\mathfrak{p}(\widetilde{\rho})$, which leads to a similarity profile $\rho^{*}$ satisfying (5.17), for instance, by following the ideas of [Váz07, Thm. 16.2]. However, the decay properties of $\rho^{*}$ are less clear, and in multiple dimensions, they may depend on the angular profile $\hat{\rho}$ and assumption (A) may be a proper restriction.
@@
    Let $\alpha\geq 0$, let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ with $\mathfrak{p}^{\prime}>0$ in (0,)(0$\infty$)( 0 , ∞ ), let $\rho^{*}=\hat{\rho}\geq 0$ be constant and $n^{*}=0$, and if $\rho^{*}=0$, let $\mathfrak{p}$ satisfy (2.14), or
@@
$\mu$ :$=\Big{\|}\frac{|R^{*}|}{2k(\rho^{*})^{\gamma}}+\frac{3|R^{*}|}{2\rho^{*}}\Big{\|}_{L^{\infty}({\mathbb{R}})},\qquad\quad K=\int_{\mathbb{R}}|R^{*}|\,\mathrm{d}y,:$
@@
where we used ${\mathcal{E}}(\tau;\rho,n\,|\,\rho^{*},n^{*})\leq\mathcal{R}$ for the first term in the integral on the right-hand side. This inequality resembles (4.2) in the proof of Theorem 2.3. Invoking now Gronwall’s inequality from Lemma 3.15 in the same way, we conclude (5.18) in this case.
@@
Theorem 5.10 yields convergence of energy-variational solutions towards the similarity profile under the same flatness condition as for weak solutions, see Theorem 2.3. In a similar way as in Section 4, we could also deduce convergence of the time-averaged kinetic energy. In Theorem 5.10, we further deduce the convergence of the energy defect to zero, which means that energy-variational solutions converge time-asymptotically to weak solutions in a certain sense. Observe that in the case $\rho_{+}=\rho_{-}$ the result holds for any regularity weight $\mathcal{K}$, while in the case $\rho_{+}\neq\rho_{-}$ we require the specific choice (5.6). Clearly, one could also take a regularity weight that is larger than (5.6), but this would lead to a more restrictive flatness condition.